Grade 9General Physics

The distance between 2nd crest and 6 trough of a wave shown below is 24 cm if the velocity of the moving wave is 20m/s then what is the frequency of the wave

Harsh
9 Years agoGrade 9

2 Answers

Vikas TU
9 Years ago
Distance b/w one crest and one trough is lemda/2 and b/w two crests or two trough it is lemda.
Therfore from 2nd crest to 6th trough the total distance = 4lemda + lemda/2 => 9lemda/2
that is given as:
9lemda/2 = 24 cm
lemda= 16/3 cm
V = w/k => w*lemda/2pi => w = 2*3.14*20*3/16 => 23.55 rad/s
Acoustic Dhruv
8 Years ago
Here,the wavelength is 4.5 lambda which is 24 cm. So the wavelength is 0.24/4.5 m.f=v/lambdaSo by putting values it becomes 375 Hz